OpenStudy (anonymous):

The graph of which of the following will be parallel to the graph of 4x – 3y = –12?

OpenStudy (anonymous):

put line 4x - 3y = -12 in slope form y = mx+c so eq will be 3y = 4x +12 y = 4/3 x +4 now the graph parallel to this will have same slope,m m = 4/3